Traditional Forecasting application (Due:09/1818) The XYZ company introduced the new technology (currently advertized as U- VERSE) in mid 2010. At the time, the company estimated that their new technology will have a long-term market share of 70% (ie, M-70%) According to the companys records, by January Ist, 201, its market share had reached 5% of its peak, and it reached to 45% of its peak by April first 2013 The companys long-term technology plan calls for starting R&D activities for new technology 2 years before the current technology reaches its peak (or within 1% of its peak). PART I: Determine 1)- The calendar year (call this X) the companys market share will reach its expected peak, and the year R&D activities for the next technology should start 2)- Use a time unit of1 quarter (0.25 year), calculate the market share for each period, from 2013 to year X+2, and graph the corresponding S-Curve. PART I When the technology was introduced, using historical data for similar products, it was determined that A)-Demand for the product in year T will increase for some time in the future and could be estimated using the following equation D(t) = 1.1(1.2010) +0.5 Yearly demand in million units, B)-The company will realize a long-term average revenue of S150 per unit. C)-Total Cost of R&D activities, promoting the technology introduction and other expenses, will be decided from the following equation Total costs in million-dollar, TC = 260+50*(T-2010.5 +0.8 Using the info 1)- How long will it take the company to recover its investment and rov nine start generating profit. (This is called Break-Even -Point) 2)- Suppose at the end of year 2018, the company sales the technology for S400 Mil to a competitor, Generate profit or loss for the company from the financial outcome of the technol year 213 to the end of 2018. What was the net ogy
Answer #1

Part I

1. The new technology was launched in mid 2010. Let us suppose that it was launched on July 1, 2010. Uptill Jan 1, 2011 there were two quarters. In 2 quarters the market share reached 5% of its peak. By April 1, 2013, it had reached 45% of its peak. From July 1, 2010 to April 1, 2013 there are 11 quarters.

In 11 quarters it reached 45% of its peak. To reach 100% of its peak it would need 100 11 × 24.42 45 quarters from June, 2010. This corresponds to July 1, 2016. So it will reach its expected peak in 2016. So the year R&D activities for the next technology should start is 2014.

Part II

1. Suppose it breaks even 'T' years from 2010. Then the demand for the product will increase according to the equation D(t) = 1.1(T-2010)2 0.5  

So in T years from 2010 its total demand would be

  1.1(12) 0.51.1(22)0.5+.1(T0.50.5   (the additional 0.5 at the end is to consider demand in year 2010)

The sum to T terms of this series is TT+127++0.5(T+1) 1.1 6   (Since sum to n terms of the squares of first n natural numbers is [n(n+1)(2n+1)/6])

Now multiply this sum by the long term average revenue of 150. This will give the total revenue after T years.

Total revenue after T years is 150*[1.1[\frac{T(T+1)(2T+1)}{6}]+0.5(T+1)]

Total cost after T years after 2010 is given by 180+50*(T-0.5)+0.8.

At the break even point total revenue must equal total cost. So equate total revenue with total cost and solve for T.

150 1.1 lrT(T+1)( 2T +1) ] + 0.5(T + 1)] = 180 + 50 6

Solving this T=0.63551.

So it will take 0.6 years for the company to break even.